# Traversable Wormholes with Spontaneous Symmetry Breaking

### v1

@@ 3 Photon sphere, Lyapunov exponent, Shadow radius, and Innermost Stable Circular Orbits
In this section, we discuss a few additional properties of this space-time metric: the radius of the photon sphere $(r_{\rm ph})$, the shadow radius ($r_{sh}$), the innermost stable circular orbits $(r_{\textrm{ISCO}})$ and the Lyapunov exponent $(\lambda)$. While a photon sphere is associated with massless particles, an ISCO is related to massive particles [48, 49, 50, 51, 52]. In particular, the existence of a photon circular orbit plays a crucial role in understanding the characteristic signatures of the quasi-normal modes and shadow radius. We parametrize the tangent vector to the worldline of any particle (massless or massive) by an affine parameter $\tau$ and write
@@
$G(r)\left(\frac{dt}{d\lambda}\right)=E\ ;\qquad\quad\left(r^{2}+a^{2}\right)\left(\frac{d\phi}{d\lambda}\right)=L,$ (27)
where, we recall that $G(r)=1+C_{1}+\frac{C_{2}}{2a^{3}}\left[ar+(a^{2}+r^{2})\tan^{-1}(r/a)\right]$ for the exact solution considerd here. Combining 26 and 27 we derive
@@
For a massless particle $\epsilon=0$. In a spherically symmetric environment, the $r$ values which can generate $V_{0}^{{}^{\prime}}(r)=0$ give the coordinate locations sufficiently close to the mass enclosed within the metric, so that photons are restricted to travel along circular geodesic orbits only. The effective potential for such a case is written as $V_{0}(r)=G(r)\left(\frac{\mathbb{L}^{2}}{r^{2}+a^{2}}\right)$. The radius of the circular photon orbit, $r_{\rm ph}$ is calculated by setting $V_{0}^{\prime}(r)=0$ [54, 55, 56, 57], leading to the condition
$2rG=(a^{2}+r^{2})G^{\prime}~{}~{}\Rightarrow~{}~{}\frac{L^{2}(C_{2}-2r_{\rm ph})}{(a^{2}+{r_{\rm ph}}^{2})^{2}}~{}~{}\Rightarrow~{}~{}r_{\rm ph}=\frac{C_{2}}{2}.$ (30)
A circular photon orbit is stable or unstable depending on the signature of $V_{0}^{\prime\prime}$ evaluated at $r_{\rm ph}$. An explicit form can be calculated as
which is always negative and confirms the existence of unstable photon orbits in this spacetime. This behavior is independent of the choice of $a$. This is demonstrated in 7 and 8 by plotting the effective potential and its’ second derivative as a function of $a$, taking $\mathbb{L}=2$.
@@
The Lyapunov exponent is determined by studying small perturbations around the radial null trajectory. We examine how the effective potential, $V_{0}$, varies with a small perturbation $\delta r$ as the radial coordinate approaches the photon orbit at $r=r_{\rm{ph}}$. This leads us to an expression for the evolution of the perturbation, given by $\delta r\sim\exp(\lambda t)$, where $\lambda$ is the Lyapunov exponent. The Lyapunov exponent quantifies the rate at which these perturbations grow or decay over time, characterizing the stability of the orbit. It can be expressed in terms of the metric coefficient using the formula [58, 59, 60, 61]
@@
This defines a direct correlation between the dynamics of the trajectory and the underlying spacetime geometry. Using the exact solution for $G(r)$ and the radius of the photon sphere, we calculate the Lyapunov exponent as
@@
We plot $\lambda$ in 9 as a function of $a$. A positive Lyapunov exponent signifies that nearby trajectories diverge over time, indicating a strong sensitivity to initial conditions.
Figure 9: Left: Plot of the Lyapunov exponent as a function of $a$, for $C_{1}=1$ and different values of $C_{2}$. Right: Plot of the Lyapunov exponent as a function of $a$, for $C_{2}=1$ and different values of $C_{1}$.
The ”shadow radius” of a compact object refers to the apparent size of the dark shape cast by the compact object against the backdrop of light, usually from a bright source behind it. A critical impact parameter can be defined as one that corresponds to the boundary between light rays that escape to infinity and those that are captured by the geometry (which in this case is either a one-way or a traversable wormhole). The shadow radius $r_{sh}$ can be calculated by finding the critical parameter associated with null geodesics [62] using the equation
$\frac{1}{r^{2}_{\textrm{sh}}}=\frac{G(r_{\rm ph})}{(r_{\rm ph}^{2}+a^{2})}~{}~{}\Rightarrow~{}~{}r_{sh}=\sqrt{\frac{2a^{3}}{2a^{3}C_{1}+C_{2}\tan^{-1}\left(\frac{C_{2}}{2a}\right)+2a}}~{}.$ (34)
Figure 10: Left: Shadow radius as a function of $a$ for $C_{1}=1$ and different values of $C_{2}$. Right: Shadow radius as a function of $a$ for $C_{2}=1$ and different values of $C_{1}$.
In 10, we plot the shadow radius as a function of $a$ for different sets of $C_{1}$ and $C_{2}$.
In order to study the geodesic orbit for massive particles, we consider timelike worldline, i.e., $\epsilon=-1$ in 29. The innermost stable circular orbit primarily corresponds to the condition that the effective potential experienced by the particle must obey $V^{\prime}_{-1}=0$ [61]. This condition can be simplified into
@@
One can solve this equation numerically, however it is better to proceed with a fixed coordinate value $r=r_{c}$ where a circular orbit is expected to form. Thereafter, it is possible to derive a required angular momentum, $L_{c}$, in terms of $r_{c}$, $a$, $C_{1}$, and $C_{2}$. In this context, the innermost circular orbit for a massive particle will simply be the coordinate value of $r_{c}$, which can minimize $L_{c}$. Since with a $\tan^{-1}\left(\frac{r_{c}}{a}\right)$ term, it is non-trivial to proceed, once again, we study the condition within a small neighborhood around $r=0$. Within this limit we approximate $\tan^{-1}\left(\frac{r}{a}\right)\simeq\frac{r}{a}$ and neglect the higher order terms of $r_{c}$ to ultimately find
$L_{c}=\sqrt{\frac{2a^{7}C_{1}r_{c}+a^{5}C_{2}}{a^{3}(C_{2}-2r_{c})}}.$ (36)
We can compare this expression with circular orbits in weak-field GR. In classical physics, the angular momentum per unit mass for a particle with angular velocity $\omega$ is $L_{c}\sim\omega r_{c}$. Kepler’s third law of planetary motion implies that $\omega^{2}\sim\frac{G_{N}m}{r_{c}}$, where $m$ is the mass of a central object. Therefore we can write that
@@
Since our solution has no Schwarzschild limit, there is no direct mass parameter involved in this metric. However, compared with the weak field limit, we note that the parametric combination of $\frac{2a^{4}C_{1}}{C_{2}}$ might have a correlation with a corresponding mass function.
@@
We conclude with a remark and reminder that the condition $\phi\frac{dV}{d\phi}\geq 0$ is usually treated as a requirement such that the self-interaction potential has a local minimum at $\phi=0$ and no local maximum within the $\phi$-range considered. This condition serves as the key to effectively rule out non-trivial scalar deformations of the Schwarzschild solution. For the present solution it can be seen in 11 that the condition is always violated irrespective of the scalar field profile under consideration.
4 Conclusion
The exact solution presented in this article represents a system of static spherically symmetric general relativity plus a massive scalar field. The scalar self interaction is Higgs type as in a linear combination of quadratic and quartic terms. It is found that for the solution to be consistent the coefficient of the quadratic term, as in the mass term $M$, must be a function of coordinates (in the form of Eq. 4). This enforces a spontaneous symmetry breaking in the system. The solution does not exhibit any singularity as checked from the curvature scalar profiles which remain finite for all possible coordinate ranges. The behavior of radial null geodesics demonstrates that the metric represents either a one-way wormhole or a two-way traversable wormhole, depending on the parameter space. We compute and discuss properties of the photon sphere, Lyapunov exponent, shadow radius and the radius of the innermost stable circular orbits that can be accommodated by this spacetime metric. It may be an interesting prospect to compare this solution with dirty black holes [63, 64, 65], associated with non-trivial stress-energy tensor or even with regular black holes [39]. This is usually done through an analysis of the derived metric tensor in a generalized tortoise coordinate and subsequent derivation of Regge-Wheeler equations for spin zero, spin one and/or spin two particles coupled to gravity. This shall be discussed in a follow-up article. The metric solution is novel as it does not reduce to the Schwarzschild metric in any limit. Moreover, it opens up an intriguing thought: to associate spontaneous symmetry breaking with a threshold of wormhole throat formation.

### v3

@@ 4 Photon sphere, Lyapunov exponent, Shadow radius, and Innermost Stable Circular Orbits
In this section, we discuss a few additional properties of the two classes of space-time metrics described here: the radius of the photon sphere $(r_{\rm ph})$, the shadow radius ($r_{sh}$), the innermost stable circular orbits $(r_{\textrm{ISCO}})$ and the Lyapunov exponent $(\lambda)$. While a photon sphere is associated with massless particles, an ISCO is related to massive particles [66, 67, 68, 69, 70]. In particular, the existence of a photon circular orbit plays a crucial role in understanding the characteristic signatures of the quasi-normal modes and shadow radius. We parametrize the tangent vector to the worldline of any particle (massless or massive) by an affine parameter $\tau$ and write
@@
$G(r)\left(\frac{dt}{d\lambda}\right)=E\ ;\qquad\quad\left(r^{2}+a^{2}\right)\left(\frac{d\phi}{d\lambda}\right)=L.$ (37)
So far the mathematical construction does not make any particular choice of $G(r)$. We first discuss the generic calculations and remind the reader that the functional forms of $G(r)$ for the two solutions are given in 14 (for a Phantom wormhole) and 33 (for a Kiselev wormhole). Combining 36 and 37 we derive
@@
For a massless particle $\epsilon=0$. We look for the $r$ values which can generate $V_{0}^{{}^{\prime}}(r)=0$. Under spherical symmetry, these points define the coordinate locations sufficiently close to the mass enclosed within the metric, where photons are restricted to travel along circular geodesic orbits. The effective potential for such a case is written as $V_{0}(r)=G(r)\left(\frac{\mathbb{L}^{2}}{r^{2}+a^{2}}\right)$. The solution $r_{\rm ph}$ of the equation $V_{0}^{\prime}(r)=0$ gives us the radius of the circular photon orbit [72, 73, 74, 75].
@@
For a Phantom wormhole 14, this leads to the condition
$2rG=(a^{2}+r^{2})G^{\prime}~{}~{}\Rightarrow~{}~{}\frac{L^{2}(C_{2}-2r_{\rm ph})}{(a^{2}+{r_{\rm ph}}^{2})^{2}}=0,~{}~{}\Rightarrow~{}~{}r_{\rm ph}=\frac{C_{2}}{2}.$ (40)
A circular photon orbit is stable or unstable depending on the signature of $V_{0}^{\prime\prime}$ evaluated at $r_{\rm ph}$. The explicit form of $V_{0}^{\prime\prime}$ evaluated at $r_{\rm ph}$ is
which is always negative and confirms the existence of unstable photon orbits for a Phanton wormhole spacetime. This behavior is independent of the choice of $a$. We demonstrate this in 11 and 12 by plotting the effective potential and its’ second derivative for different values of $a$, taking $L=2$. Similarly, the photon orbit radius for the Kiselev wormhole metric can be calculated as
$r_{\rm ph}=-a^{3}C_{2}p,$ (42)
which indicates that a physical photon orbit can only develop if either $C_{2}$ or $p$ is chosen as negative. The second derivative of the effective potential $V_{0}^{\prime\prime}$ is evaluated at $r_{\rm ph}$ as
$V_{0}^{\prime\prime}(r)\Big{|}_{r_{\rm ph}}=-\frac{2L^{2}\left(a^{6}C_{2}^{2}p^{2}+a^{2}\right)}{\left(a^{6}C_{2}^{2}p^{2}+a^{2}\right)^{3}},$ (43)
which is, again, always negative. This confirms the existence of unstable photon orbits for the Kiselev wormhole spacetime as well. The overall behavior is independent of the choice of $a$ which we demonstrate in 13 and 14. We plot the effective potential and its’ second derivative for different values of $a$, taking $L=2$.
A Lyapunov exponent is connected to small perturbations around the radial null trajectory defined in a metric tensor of a compact object, in this case, the wormhole metrics. We discuss how the effective potential, $V_{0}$, defined in 39, is affected by a small perturbation $\delta r$ close to the photon orbit radius $r=r_{\rm{ph}}$. We derive to an expression for the evolution of the perturbation, given by $\delta r\sim\exp(\lambda t)$, where $\lambda$ is the Lyapunov exponent. This exponent quantifies the rate at which perturbations can grow or decay over time, characterizing the stability of the orbit as well as the object [76, 77, 78, 79]. The generic expression is found as
@@
Figure 15: Left: Plot of the Lyapunov exponent for a Phantom wormhole as a function of $a$, for $C_{1}=1$ and different values of $C_{2}$. Right: Plot of the Lyapunov exponent as a function of $a$, for $C_{2}=1$ and different values of $C_{1}$.
The exponent is dependent on metric coefficients, which indicate a correlation of the trajectory with the underlying spacetime geometry. Using 14 and the radius of the photon sphere for Phantom wormhole, we calculate the Lyapunov exponent as
@@
For a Phantom wormhole we plot $\lambda$ in 15 for different values of $a$. Similarly, the Lyapunov exponent for the Kiselev wormhole is shown in 16. For both of the cases, a positive Lyapunov exponent signifies that nearby trajectories diverge over time, indicating a strong sensitivity to initial conditions.
Figure 16: Left: Plot of the Lyapunov exponent for a Kiselev wormhole as a function of $a$, for $C_{1}=1$ and different values of $C_{2}$. Right: Plot of the Lyapunov exponent as a function of $a$, for $C_{2}=1$ and different values of $C_{1}$.
The shadow radius for a compact object refers to the apparent size of the dark shape cast by the object against a backdrop of light, usually from a bright source behind it. One can define a critical impact parameter which corresponds to the boundary between light rays that escape to infinity and those that are captured by the geometry (which in this case is either a one-way or a traversable wormhole). For a Phantom wormhole, the shadow radius $r_{sh}$ is calculated by finding the critical parameter associated with null geodesics [80] for 14,
$\frac{1}{r^{2}_{\textrm{sh}}}=\frac{G(r_{\rm ph})}{(r_{\rm ph}^{2}+a^{2})}~{}~{}\Rightarrow~{}~{}r_{sh}=\sqrt{\frac{2a^{3}}{2a^{3}C_{1}+C_{2}\tan^{-1}\left(\frac{C_{2}}{2a}\right)+2a}}~{},$ (46)
Figure 17: Shadow radius of a Phantom wormhole for different values of $a$. Left: $C_{1}$ is fixed at $1$ while different values of $C_{2}$ are chosen. Right : $C_{2}$ is fixed at $1$ while different values of $C_{1}$ is considered.
We plot the shadow radius in 17 as a function of $a$ for different sets of $C_{1}$ and $C_{2}$. Similarly, the shadow radius for the Kiselev wormhole is found as
$r_{sh}=\sqrt{\frac{a^{3}}{a^{3}C_{1}p+a^{3}C_{2}p\tan^{-1}\left(a^{2}C_{2}p\right)+a+m}}~{},$ (47)
which is plotted in 18.
Figure 18: Shadow radius of a Kiselev wormhole for different values of $a$ for Left : $C_{1}=1$ and different values of $C_{2}$ ; Right : $C_{2}=1$ and different values of $C_{1}$.
Geodesic orbits for massive particles can be described using a timelike worldline for the $\epsilon=-1$ case of 39. The innermost stable circular orbit then corresponds to the condition that the effective potential experienced by the particle must obey $V^{\prime}_{-1}=0$ [79]. For the Phantom wormhole, this can be simplified into
@@
One can solve this equation numerically, however it is better to proceed with a fixed coordinate value $r=r_{c}$ where a circular orbit is expected to form. At this coordinate value we derive the required angular momentum $L_{c}$, in terms of $r_{c}$, $a$, $C_{1}$, and $C_{2}$. Therefore the innermost circular orbit for a massive particle forms at the coordinate value $r_{c}$ for which $L_{c}$ is minimized. Since with a $\tan^{-1}\left(\frac{r_{c}}{a}\right)$ term, it is non-trivial to proceed, we study the condition within a small neighborhood around $r=0$. In this limit we approximate $\tan^{-1}\left(\frac{r}{a}\right)\simeq\frac{r}{a}$ and neglect the higher order terms of $r_{c}$ to write
$L_{c}=\sqrt{\frac{2a^{7}C_{1}r_{c}+a^{5}C_{2}}{a^{3}(2r_{c}-C_{2})}}.$ (49)
This expression can be compared with circular orbits in weak-field GR. In classical physics, the angular momentum per unit mass for a particle with angular velocity $\omega$ is $L_{c}\sim\omega r_{c}$. Kepler’s third law of planetary motion implies that $\omega^{2}\sim\frac{G_{N}m}{r_{c}}$, where $m$ is the mass of a central object. Therefore one can write that
@@
In comparison with the weak field limit, we note that for a Phantom wormhole, the parametric combination of $\frac{2a^{4}C_{1}}{C_{2}}$ has a connection with the corresponding mass function. Similarly, the $L_{c}$ for a Kiselev wormhole is derived as
$L_{c}=\frac{\sqrt{-a^{5}C_{2}p+a^{4}C_{1}pr_{c}+amr_{c}}}{\sqrt{a^{3}C_{2}p+r_{c}}}.$ (51)
5 Conclusion
The idea of matter holds particular importance in GR. It is the connection between an Einstein tensor and an energy-momentum tensor (both being free of covariant divergence) that fabricates GR out of Riemannian geometry. It can not be denied that the field equations describing this connection between geometry and matter, can produce exotic solutions. Some of these solutions carry space-time points of infinite curvature within their boundary/horizon while some them behave a bit more radically and allow a chance of safe passage through to a distant part of the universe. The genesis of the matter and energy that allows a formation of such exotic objects remains an interesting topic of research. To that end, the present article explores three notions : (i) it checks if a non-trivial scalar deformation of the Schwarzschild solution can exist for a static spherically symmetric GR plus scalar field system ; (ii) it explores the method of regularization of an existing singular solution of GR and the properties of the resulting geometry and (iii) it forms a connection between the formation of wormhole throat to an effective phase transition of the matter field involved, through a spontaneous symmetry breaking.
@@ 5 Conclusion
For any massive (self-interaction potential $V(\phi)$) scalar field minimally coupled to GR, the condition $\phi\frac{dV}{d\phi}\geq 0$ is usually treated as a requirement such that the self-interaction potential has a local minimum at $\phi=0$ and no local maximum within the $\phi$-range considered. This condition can be treated as a key to effectively ruling out non-trivial scalar deformations of the Schwarzschild solution. For the present solution it can be seen in 19 that the condition is always violated irrespective of the scalar field profile under consideration.
The solutions we explore in this article describe two-way traversable wormholes or, in selected cases, regular black holes. The first solution can be traced back to the regular black hole solution of Bronnikov [39], which is supported by a phantom scalar field (negative kinetic energy). We connect the metric with a second class of solution popular as the Kiselev black hole [40]. We derive a generalized version of the Kiselev metric. We prove that both of these solutions can be supported by a massive scalar field whose self-interaction is Higgs type, i.e., has a linear combination of quadratic and quartic terms. Moreover, the solutions are consistent if the coefficient of the quadratic term, as in the mass term $M$, is a function of coordinates (in the form of 4). The functional form of $M$ is solved from the field equations and it can be seen that the profile carries a signature of the breakdown of $Z_{2}$ symmetry in the system, around the wormhole throat radius. The solutions do not exhibit any singularity as checked from the curvature scalar profiles which remain finite for all possible coordinate ranges. The behavior of radial null geodesics demonstrates that the metrics represent either of the three: a one-way wormhole, a two-way traversable wormhole or a regular blackhole, depending on the parameter space. We compute and discuss properties of the photon sphere, Lyapunov exponent, shadow radius and the radius of the innermost stable circular orbits for the metrics. It may be an interesting prospect to further generalize or compare these solutions with dirty black holes [81, 82, 83], associated with non-trivial stress-energy tensors [57]. This is usually done through the analysis of a derived metric tensor in generalized tortoise coordinates. A subsequent derivation of the Regge-Wheeler equations follows for spin zero, spin one, and/or spin two particles coupled to gravity. This will be discussed in a follow-up article.